40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Sep 20 12:20:31 2021, Paco, Summary, SUS, PRM and BS Angular Actuation transfer function magnitude measurements PRM_PIT_ACT_TF.pdfPRM_YAW_ACT_TF.pdfBS_PIT_ACT_TF.pdfBS_YAW_ACT_TF.pdf
    Reply  Mon Sep 20 14:22:00 2021, rana, Summary, SUS, PRM and BS Angular Actuation transfer function magnitude measurements 
       Reply  Thu Sep 23 15:29:11 2021, Paco, Summary, SUS, PRM and BS Angular Actuation transfer function magnitude measurements BS_PRM_ANG_ACT_TF.pdf
          Reply  Fri Oct 1 14:25:27 2021, yehonathan, Summary, SUS, PRM and BS Angular Actuation transfer function magnitude measurements MICH_PRM_-0.34.pngMICH_PRM_-0.1.png
          Reply  Mon Oct 4 16:00:57 2021, Yehonathan, Summary, SUS, PRM and BS Angular Actuation transfer function magnitude measurements BS_PRM_ANG_ACT_TF_20211004.pdfBS_out_mat_20211004.txtPRM_out_mat_20211004.txt
             Reply  Mon Oct 4 16:10:09 2021, rana, Summary, SUS, PRM and BS Angular Actuation transfer function magnitude measurements 
             Reply  Tue Oct 5 20:04:22 2021, Paco, Summary, SUS, PRM and BS Angular Actuation transfer function magnitude measurements 
                Reply  Wed Oct 6 15:39:29 2021, Anchal, Summary, SUS, PRM and BS Angular Actuation transfer function magnitude measurements 
             Reply  Tue Oct 12 11:32:54 2021, Yehonathan, Summary, SUS, PRM and BS Angular Actuation transfer function magnitude measurements modified_output_matrices_radar_plots.pngnormal_output_matrices_radar_plots.png
                Reply  Tue Oct 12 16:39:52 2021, rana, Summary, SUS, PRM and BS Angular Actuation transfer function magnitude measurements 
                   Reply  Thu Oct 14 13:40:49 2021, Yehonathan, Summary, SUS, PRM and BS Angular Actuation transfer function magnitude measurements BSPRM_Actuation_Radar_plots.pngITM_Actuation_Radar_plots.png
Message ID: 16383     Entry time: Tue Oct 5 20:04:22 2021     In reply to: 16374     Reply to this: 16385
Author: Paco 
Type: Summary 
Category: SUS 
Subject: PRM and BS Angular Actuation transfer function magnitude measurements 

[Paco, Rana]

We had a look at the BS actuation. Along the way we created a couple of issues that we fixed. A summary is below.

  1. First, we locked MICH. While doing this, we used the /users/Templates/ndscope/LSC/MICH.yml ndscope template to monitor some channels. I edited the yaml file to look at C1:LSC-ASDC_OUT_DQ instead of the REFL_DC. Rana pointed out that the C1:LSC-MICH_OUT_DQ (MICH control point) had a big range (~ 5000 counts rms) and this should not be like that.
  2. We tried to investigate the aforementioned thing by looking at the whitening / uwhitening filters but all the slow epics channels where "white" on the medm screen. Looking under CDS/slow channel monitors, we realized that both c1iscaux and c1auxey were weird, so we tried telnet to c1iscaux without success. Therefore, we followed the recommended wiki procedure of hard rebooting this machine. While inside the lab and looking for this machine, we touched things around the 'rfpd' rack and once we were back in the control room, we couldn't see any light on the AS port camera. But the whitening filter medm screens were back up.
  3. While rana ssh'd into c1auxey to investigate about its status, and burtrestored the c1iscaux channels, we looked at trends to figure out if anything had changed (for example TT1 or TT2) but this wasn't the case. We decided to go back inside to check the actual REFL beams and noticed it was grossly misaligned (clipping)... so we blamed it on the TTs and again, went around and moved some stuff around the 'rfpd' rack. We didn't really connect or disconnect anything, but once we were back in the control room, light was coming from the AS port again. This is a weird mystery and we should systematically try to repeat this and fix the actual issue.
  4. We restored the MICH, and returned to BS actuation problems. Here, we essentially devised a scheme to inject noise at 310.97 Hz and 313.74. The choice is twofold, first it lies outside the MICH loop UGF (~150 Hz), and second, it matches the sensing matrix OSC frequencies, so it's more appropriate for a comparison.
  5. We injected two lines using the BS SUS LOCKIN1 and LOCKIN2 oscilators so we can probe two coils at once, with the LSC loop closed, and read back using the C1:LSC-MICH_IN1_DQ channel. We excited with an amplitude of 1234.0 counts and 1254 counts respectively (to match the ~ 2 % difference in frequency) and noted that the magnitude response in UR was 10% larger than UL, LL, and LR which were close to each other at the 2% level.

[Paco]

After rana left, I did a second pass at the BS actuation. I took TF measurements at the oscilator frequencies noted above using diaggui, and summarize the results below:

TF UL (310.97 Hz) UR (313.74 Hz) LL (310.97 Hz) LR (313.74 Hz)
Magnitude (dB) 93.20 92.20 94.27 93.85
Phase (deg) -128.3 -127.9 -128.4 -127.5

This procedure should be done with PRM as well and using the PRCL instead of MICH.

ELOG V3.1.3-